OpenStudy (rational):

how to solve recurrence relations with exponents solve \(\large a_n = 1-2a_{n-1}^2~;~a_0 = -\frac{1}{3}\)

OpenStudy (anonymous):

it's always useful to try substitution until you see a patern

OpenStudy (rational):

like this ? \[\large \begin{align} \\ a_n &= 1-2a_{n-1}^2 \\~\\&= 1 - 2(1-2a_{n-2}^2)^2\\~\\\end{align}\]

OpenStudy (anonymous):

exactly. it's going to be very tedious because of the square exponent

OpenStudy (ikram002p):

why is difficult ?

OpenStudy (ikram002p):

its*

OpenStudy (rational):

can this be a solution : \[\large a_n = -\cos(2^n \arccos(-1/3)) \] even if this is a solution, this doesn't look that exciting... im feeling there should be a way to find solution without trig -.-

OpenStudy (ikram002p):

how did u got that ?

OpenStudy (rational):

from double angle identity \[\cos (2\theta) = 2\cos^2\theta - 1\]

OpenStudy (ikram002p):

u know we could always check revercly , then find limit when n goes to \(\infty\) that would be extra solution